OpenStudy (anonymous):

Jen is building an igloo. (Hint: the shape is half a sphere). What radius should she make the igloo in order for it to have a volume of 20 ft ^3? round answer to 2 decimal places

OpenStudy (hoblos):

the volume of the igloo is \[\frac{ 4 }{ 3\times2}\pi r ^{3} = 20\] \[r ^{3} = \frac{ 20\times6 }{ 4\pi} = \frac{ 30 }{ \pi }\] \[r = \sqrt[3]{\frac{ 30 }{ \pi }} \]